It would be nice if someone could show me how to solve this mixture problem. Thanks in advance.
A 15 liter bucket of soda contains a mixture with a concentration of 50% corn syrup. How much of this mixture x must be withdrawn and replaced by 100% corn syrup to bring the mixture up to a concentration of 60% corn syrup?
x=________liters

Answer by Paul(988)

100x+50(15-x)=60(15)
100x+750-50x=900
Simplify
50x=150 (divide both sides 50)
x=3
Hence for 100% it should bee 3L and for 50% it should be 12L